OpenStudy (anonymous):

tan40 +tan 135+tan 120

OpenStudy (paxpolaris):

tan (135) = tan (180 - 45) = - tan (45)

OpenStudy (paxpolaris):

tan (120 ) = tan (180-60) = - tan60

OpenStudy (anonymous):

=zero